Solve the given integral equation or integro-differential equation for y(t). y(t+ft-vy(v) dv=t, y(0)=0 0 y(t) =
Answer #1

+ Hol=0 y (t) + s (t-v) yu) do = t , differentiating both side, | ! ***-eywo] - - y (A) + d at . ) Leibnitz Rule: Let F(x,=) (MH) (m²m+1)-0 of mat, me 1 + 3 i (t)= Get + 2 ( le cose + + Cyline sin St y (t)= - Get & Cocos B tot el 12+ g. (-ain. )by , y(o)=1 using y (o)=1 in equation 1 - 9+9+9 = 1=3G 9 = 1² 3 now, G=% a = -6= = // (z=634= therefore by W. #2 y(t) = $ e